# Exceptional topology in Non-Hermitian Twisted Bilayer Graphene

### v2

@@ Abstract
Twisted bilayer graphene has extraordinary electronic properties at the magic angle along with an isolated flat band at magic angle. However, the non-Hermitian phenomena in twisted bilayer graphene remain unexplored. In this work, we study a non-Hermitian TBG formed by one-layer graphene twisted relative to another layer with gain and loss. Using a non-Hermitian generalization of Bistritzer-MacDonald model, we find Dirac cones centered at only $K_{M}$ ($K^{\prime}_{M}$) corner of the moiré Brillouin zone at $K^{\prime}$ ($K$) valley deformed in the presence of non-Hermiticity. This is different from single layer graphene with gain and loss, where rings of exceptional points appear in both $K$ and $K^{\prime}$ corners of the Brillouin zone. The coincident of exceptional rings at $\Gamma_{M}$ point characterizes an “exceptional magic angle”, at which the system hosts flat bands with zero energy and finite lifetime. More interestingly, we find that the topological charge in the moiré Brillouin zone is conserved during the expansion and fusion of the exceptional ring, which is absent in two-dimensional systems constraining by Nielsen-Ninomiya theorem. These findings can be demonstrated in realistic cold atom and metamaterial systems and will stimulate further study on non-Hermitian phenomena in twistronic.
@@
Figure 1: (color online)Schematics of twisted bilayer graphene with onsite gain and loss. (a) Real-space geometrical picture: The gray (orange) honeycomb lattice marks the top (bottom) layer. The red (blue) dots represent the amplifying (lossy) sites in the bottom layer. (b) Momentum-space geometrical picture: The gray (brown) lattice marks the Brillouin zone of the top (bottom) layer graphene. Their expansion around the $K$ ($K^{\prime}$) valley forms moiré Brillouin zone in green (blue). (c) Moiré Brillouin zone physics at $K^{\prime}$ valley: Under the effect of onsite gain and loss, Dirac points at $K_{M}$ corners (related to $K_{b}$ points of bottom layer graphene) morph into exceptional rings. At an “exceptional magic angle”, the exceptional rings touch at $\Gamma_{K}$ point. (d) Moiré Brillouin zone physics at $K$ valley: exceptional rings from Dirac points at $K^{\prime}_{M}$.
@@
Graphene owes extraordinary electronic and optical properties in two-dimensional (2D) systems  [1, 2]. These properties relate to its lattice structure, which is directly interconnected with its topological property. In single-layer graphene, the presence of unusual linear band dispersion termed “Dirac cone” at $K$ and $K^{\prime}$ points of the Brillouin zone is derived from its honeycomb lattice geometry. The two Dirac points are related by time-reversal symmetry and thus have opposite Berry phases, which has been confirmed in experiments [3, 4].
@@
When two graphene layers are shifted with each other in rational competing periodicities upon shearing or twisting, they form a moiré pattern [5, 6]. One prominent example is a graphene bilayer with a relatively small angle rotation between the layers [7]. The recent studies on twisted bilayer graphene (TBG) focus on the extraordinary electronic properties at the magic angle along with an isolated flat band [8, 9, 10, 11, 12]. It is related to a variety of exotic phases including quantum anomalous Hall [13, 14, 15, 16, 17], fractional Chern insulators [18, 19, 20], ferromagnetic states [14, 21, 22], nematicity [23], unconventional superconductivity [24, 25, 26, 27, 28, 29]. On the other hand, twisted bilayer graphene has extraordinary topological properties. Since the two Dirac points in moiré Brillouin zone (mBZ) in one valley emanate from different layers, their Berry phases can be identical due to the symmetry interlayer hopping term, which is generally not allowed in two-dimensional periodic systems following the Nielsen-Ninomiya theorem [30, 31]. This is in contrast to single-layer graphene, in which the Berry phases of Dirac cones in different valleys has different topological charges.
@@
The topological properties of two-dimensional systems can be generalized to non-Hermitian systems by introducing on-site gain and loss [32, 33, 34]. The presence of non-Hermiticity can transform Hermitian degenerate points, such as a Dirac-like point or Weyl point, into a ring of exceptional points [35, 36]. At an exceptional point, not only the real and imaginary parts of the eigenvalues degenerate, but also the eigenvectors coalesce. Many phenomena are unique to non-Hermitian systems, including non-Hermitian skin effect [37] and non-Hermitian topological classifications [38, 39]. Despite topological robustness has been shown in one-dimensional moiré lattice under stain, the non-Hermitian effect of twisted moiré system with magic angle has been overlooked [40, 41].
@@
Usually, non-Hermiticity is detrimental to the topology in single-layer graphene. When the exceptional contours with opposite topological charges merge, the topological charge can be dissipated [36]. In single-layer graphene, the annihilation of exceptional points from different valleys has been experimentally observed in a two-dimensional photonic system [42]. Since the moire BZ valley of TBG has different topological properties from single-layer graphene, it is natural to ask how the exceptional geometry appears and develops in a non-Hermitian TBG system.
@@
In this paper, we uncover a new type of twisted band engineering in TBG, with non-Hermitian perturbation on one of the layers. As shown in Fig. 1(a), balanced gain and loss are put on the bottom layers (in red and blue, respectively). The superposing of two graphene layers with a twist angle between them creates moiré patterns. In the mBZ shown in Fig. 1(b), the low-energy band structure consists of Dirac cones from the rotated layers located at the $K_{M}$ and $K^{\prime}_{M}$ corners. When the non-Hermitian perturbation is turned on the bottom layer, Dirac points at $K_{M}$ ($K^{\prime}_{M}$) corner in the mBZ at $K^{\prime}$ ($K$) valley morph into exceptional rings and the corresponding Dirac cones become exceptional cones, as shown in Fig. 1(c) and (d) respectively. We find that this non-Hermitian system hosts special angles for which multiple exceptional cones coincide at $\Gamma_{M}$ points in an mBZ. We call this phenomenon “exceptional magic”. And importantly, our results demonstrated that the total Berry charge is conserved in the mBZ even after the merging of exceptional rings, this is a large difference from non-Hermitian single-layer graphene.
@@
The paper is organized as follows. Section II introduces the non-Hermitian TBG model. Both the real and imaginary parts of the low-energy moiré band spectrum will be explored in Sec. III. In Sec. IV, we will use a non-Hermitian three triod model to discuss exceptional physics. In Sec. V, we calculate Beery curvature and the corresponding Chern number. Finally, we conclude our results in Sec. VI.
@@ II The non-Hermitian bilayer model
We consider balanced gain and loss on the bottom layer of the bilayer system. We can write down a non-Hermitian generalization of the single-valley model for twisted bilayer graphene. The low-energy Hamiltonian for the mBZ at $K$ or $K^{\prime}$ valley is:
$H=\begin{pmatrix}h_{b}(\bf{k})&\bf{T}(\bf{r})\\ \bf{T}(\bf{r})^{\dagger}&h_{t}(\bf{k})\end{pmatrix}$ (1)
@@
$h_{t}=-iv_{F}\bm{\sigma}_{\theta/2}\bm{\nabla}$ (2)
and the bottom layer Hamiltonian is non-Hermitian
$h_{b}=-iv_{F}\bm{\sigma}_{-\theta/2}\bm{\nabla}+i\lambda_{V}\sigma_{z}.$ (3)
Here, $i\lambda_{V}\sigma_{z}$ is a balanced gain and loss imposing on the bottom layer, which induces non-Hermiticity. And $\bm{\sigma}_{\theta/2}$ corresponds to the rotated Pauli matrices $e^{-i\theta\sigma_{z}/4}(\pm\sigma_{x},\sigma_{y})e^{i\theta\sigma_{z}/4}$ with $\pm$ for $K$ and $K^{\prime}$ valley respectively. $\textbf{T}(\textbf{r})=\sum^{3}_{j=1}T_{j}e^{-ik_{\theta}\textbf{q}_{j}\cdot\textbf{r}}$ is the sublattice-dependent moiré potential that couples the two layers. The interlayer coupling is in the form
$T_{j+1}=w_{0}\sigma+w_{1}(\sigma_{x}\cos j\phi+\sigma_{y}\sin j\phi)$ (4)
with $\phi=2\pi/3$ and $w_{0}(w_{1})$ the interlayer coupling strength in the AA (AB) region. The unit vectors $\textbf{q}_{1}=(0,-1)$, $\textbf{q}_{2,3}=(\pm\sqrt{3}/2,1/2)$. $k_{\theta}=2k_{D}\sin(\theta/2)$ is the moiré modulation vector and $k_{D}=4\pi/(3a_{0})$ is the magnitude of the Dirac wave vector, where $a_{0}$ is the lattice constant of monolayer graphene. The hopping strength $w$ encodes the inter-layer coupling in the AB region.
@@
If there is no AA stacking ($w_{0}=0$), the Hamiltonian preserves chiral symmetry $\mathcal{C}\mathcal{H}(\textbf{r})\mathcal{C}^{\dagger}=-\mathcal{H}(\textbf{r})$ under chiral symmetry operator $\mathcal{C}=\sigma_{z}$ in the absence of non-Hermiticity [44, 45, 46, 43].
In the presence of the balanced gain and loss $i\sigma_{z}$ term, the chiral symmetry is broken but non-Hermitian chiral symmetry is preserved In the following numerical calculation, we take $\hbar v_{F}/a=2380meV$, $w_{1}$ = 110meV and $2v_{0}k_{D}=19.81eV$. They correspond to the first magic angle $\alpha\approx 0.586$.
III The appearance and development of exceptional rings
Figure 2: (Color online) (a) The real part and (b) the imaginary part of the band spectra in the mBZ for $\theta=1,u_{0}=0,u_{1}=0.005,\lambda_{V}=0.05$.
Figure 3: (color online) (a) The real part and (b) the imaginary part of the four lowest bands in the mBZ for $u_{0}=0,u_{1}=0.005,\lambda_{V}=0.072$. The real parts of the two lowest bands are flat.
@@ III The appearance and development of exceptional rings
Let us investigate the band structure for twist angle $\theta=1$. In the absence of non-Hermiticity, this is not a magic angle with a flat band. If the interlayer coupling effect is absent ($w_{0}=w_{1}=0$), the dispersion in the mBZ is equivalent to the folded dispersion in the Brillouin zone of graphene. Since the physics in $K$ valley and $K^{\prime}$ valley are related, we only discuss the $K$ valley in the following. The Dirac point at the $K$ corner in the top layer is mapped to $K_{M}$ point in the mBZ, while that at the $K$ corner of the bottom layer is mapped to $K^{\prime}_{M}$ point in the mBZ.
@@
In the presence of non-Hermitian perturbation $\lambda_{V}$, the energy dispersion becomes complex. In Fig. 2(a), we can see that non-Hermitian perturbation deforms the Dirac cones at $K^{\prime}_{M}$ corners. More specifically, the degeneracy point at the $K^{\prime}_{M}$ points in the real part of the spectrum morphs into a ring of degeneracy points. This is a stark difference from single-layer graphene with onsite gain and loss. The latter has deformed Dirac cones in both $K$ and $K^{\prime}$ corners in the Brillouin zone. This difference is due to the fact that the Dirac cones at $K_{M}$ and $K_{M}^{\prime}$ point in non-Hermitian TBG are folded from different graphene layers. The non-Hermitian $\lambda_{V}$ can only deform the Dirac cone on the bottom layer.
@@
In particular, the degeneracy of the two lowest energy states leads to the appearance of singularity points in momentum space. At the same time, a ring of degeneracy point appears at the same momentum position in the imaginary part of the spectrum, as shown in Fig. 2(b). We found that not only the eigenvalues coalesce but also their corresponding eigenvectors coalesce at this point. This confirms that these singularity points are exceptional points in momentum space. Thus, the Dirac cone is turned into a ring of exceptional points, this is similar to that in non-Hermitian semimetals [35].
At a critical non-Hermitian strength $\lambda_{Vc}$, the exceptional point approach $\Gamma_{M}$ point, where the real part whole lowest band becomes flat, as shown in Fig. 3. This characterizes an “exceptional magic” angle. According to the geometry relationship, $\lambda_{Vc}$ is determined by the lattice size of mBZ $k_{\theta}$ and Fermi velocity $v_{F}$.
IV Exceptional cone

### v3

@@ Abstract
Twisted bilayer graphene (TBG) has extraordinary electronic properties at the magic angle along with an isolated flat band at the magic angle. However, the non-Hermitian phenomena in twisted bilayer graphene remain unexplored. In this work, we study a non-Hermitian TBG formed by one-layer graphene twisted relative to another layer with gain and loss. Using a non-Hermitian generalization of the Bistritzer-MacDonald model, we find Dirac cones centered at only the $K_{M}$ ($K^{\prime}_{M}$) corner of the moiré Brillouin zone at the $K^{\prime}$ ($K$) valley deform into rings of exceptional points in the presence of non-Hermiticity, which is different from single-layer graphene with gain and loss, where exceptional rings appear in both $K$ and $K^{\prime}$ corners of the Brillouin zone. We show that the exceptional rings are protected by non-Hermitian chiral symmetry. More interestingly, at an “exceptional magic angle” larger than the Hermitian magic angle, the exceptional rings coincide and form non-Hermitian flat bands with zero energy and a finite lifetime. These non-Hermitian flat bands in the moiré system, which are isolated from dispersive bands, are distinguished from those in non-Hermitian frustrated lattices. In addition, we find that the non-Hermitian flat band has topological charge conserved in the moiré Brillouin zone, which is allowed for analogs of non-Hermitian fractional quantum Hall states.
Figure 1: (color online)Schematics of twisted bilayer graphene with onsite gain and loss. (a) Real-space geometrical picture. The gray (orange) honeycomb lattice marks the top (bottom) layer. The red (blue) dots represent the amplifying (lossy) sites in the bottom layer. (b) Momentum-space geometrical picture. The gray (brown) lattice marks the Brillouin zone of the top (bottom) layer graphene. Their expansion around the $K$ ($K^{\prime}$) valley forms the moiré Brillouin zone in green (blue). (c) Moiré Brillouin zone physics at the $K^{\prime}$ valley. Under the effect of on-site gain and loss, Dirac points at the $K_{M}$ corners (related to $K_{b}$ points of bottom layer graphene) morph into exceptional rings. At an “exceptional magic angle”, the exceptional rings touch at the $\Gamma_{K}$ point. (d) Moiré Brillouin zone physics at the $K$ valley. Exceptional rings from Dirac points at $K^{\prime}_{M}$ are shown.
@@
Graphene has extraordinary electronic and optical properties in two-dimensional systems  [1, 2]. These properties relate to its lattice structure, which is directly interconnected with its topological property. In single-layer graphene, the presence of unusual linear band dispersion termed the “Dirac cone” at $K$ and $K^{\prime}$ points of the Brillouin zone is derived from its honeycomb lattice geometry. The two Dirac points are related by time-reversal symmetry and thus have opposite Berry phases, which has been confirmed in experiments [3, 4].
@@
When two graphene layers are shifted with each other in rational competing periodicities upon shearing or twisting, they form a moiré pattern [5, 6]. One prominent example is a graphene bilayer with a relatively small-angle rotation between the layers [7]. The recent studies on twisted bilayer graphene (TBG) have focused on the extraordinary electronic properties at the magic angle along with an isolated flat band [8, 9, 10, 11, 12]. It is related to a variety of exotic phases including the quantum anomalous Hall effect [13, 14, 15, 16, 17], fractional Chern insulators [18, 19, 20], ferromagnetic states [14, 21, 22], nematicity [23], and unconventional superconductivity [24, 25, 26, 27, 28, 29]. On the other hand, twisted bilayer graphene has extraordinary topological properties. Since the two Dirac points in moiré Brillouin zone (mBZ) in one valley emanate from different layers, their Berry phases can be identical due to the symmetry interlayer hopping term, which is generally not allowed in two-dimensional periodic systems following the Nielsen-Ninomiya theorem [30, 31]. This is in contrast to single-layer graphene, in which the Berry phases of Dirac cones in opposite valleys have different topological charges.
@@
The topological properties of two-dimensional systems can be generalized to non-Hermitian systems by introducing on-site gain and loss [32, 33, 34]. The presence of non-Hermiticity can transform Hermitian degenerate points, such as a Dirac-like point or a Weyl point, into a ring of exceptional points [35, 36]. At an exceptional point, not only do the real and imaginary parts of the eigenvalues degenerate, but also the eigenvectors coalesce. Many phenomena are unique to non-Hermitian systems, including the non-Hermitian skin effect [37] and non-Hermitian topological classifications [38, 39]. Although topological robustness has been shown in one-dimensional moiré lattice under stain, the non-Hermitian effect of the twisted moiré system with magic angles has been overlooked [40, 41].
One important question is whether the existence of flat bands is compatible with non-Hermiticity. Non-Hermitian flat bands have been investigated in non-Hermitian Lieb lattices, kagome lattices, and other non-Hermitian geometrical frustrated lattices [42, 43, 44, 45, 46], which are protected by either $\mathcal{P}\mathcal{T}$ symmetry or non-Hermitian chiral symmetry [42, 44, 47]. However, these non-Hermitian flat bands in simple geometrical frustrated lattices are usually embedded in dispersive bands. From the fact that frustration between Hermitian and non-Hermitian kinetic energy in simple lattices can induce non-Hermitian flat bands, how the non-Hermitian dispersive energy scale interplays with the interlayer tunneling energy scale modulated by the moiré pattern becomes interesting.
@@
Another important question is whether the exceptional geometry in mBZ can be topological. Usually, non-Hermiticity is detrimental to the topology in single-layer graphene. When the exceptional contours with opposite topological charges merge, the topological charge can be dissipated [36]. In single-layer graphene, the annihilation of exceptional points from different valleys has been experimentally observed in a two-dimensional photonic system [48]. Since the mBZ valley of TBG has topological properties different from those of single-layer graphene, topology is not natural in non-Hermitian TBG systems.
@@
To answer these two questions, we add non-Hermitian perturbation on one of the twisted layers. As shown in Fig. 1(a), balanced gain and loss are put on the bottom layer (in red and blue, respectively). The superposing of two graphene layers with a twist angle between them creates moiré patterns. In the mBZ shown in Fig. 1(b), the low-energy band structure consists of Dirac cones from the rotated layer located at the $K_{M}$ and $K^{\prime}_{M}$ corners. When the non-Hermitian perturbation is turned on the bottom layer, Dirac points at the $K_{M}$ ($K^{\prime}_{M}$) corner in the mBZ at the $K^{\prime}$ ($K$) valley morph into exceptional rings and the corresponding Dirac cones become exceptional cones, as shown in Figs. 1(c) and 1(d) respectively. We find that this non-Hermitian system hosts special angles for which multiple exceptional cones coincide at $\Gamma_{M}$ points in an mBZ. We call this phenomenon “exceptional magic”. And importantly, our results demonstrated that the total Berry charge is conserved in the mBZ even after the merging of exceptional rings; this is a large difference from non-Hermitian single-layer graphene.
@@
The paper is organized as follows. Section II introduces the non-Hermitian TBG model. Section III shows the presence of exceptional rings in the real and imaginary parts of the low-energy moiré band spectrum. In Sec. IV, we use a non-Hermitian three-tripod model to show that the non-Hermitian chiral symmetry protects the exceptional rings. In Sec. V, we will show the existence of a non-Hermitian flat band at an exceptional magic angle larger than the Hermitian magic angle. In Sec. VI, we calculate the Berry curvature and the corresponding Chern number. Finally, we conclude our results in Sec. VII.
@@ II The non-Hermitian bilayer model
We consider balanced gain and loss on the bottom layer of the bilayer system. We can write down a non-Hermitian generalization of the single-valley model for twisted bilayer graphene. The low-energy Hamiltonian for the mBZ at the $K$ or $K^{\prime}$ valley is
$H=\begin{pmatrix}h_{b}(\bf{k})&\bf{T}(\bf{r})\\ \bf{T}^{\dagger}(\bf{r})&h_{t}(\bf{k})\end{pmatrix},$ (1)
@@
$h_{t}=-i\hbar v_{F}\bm{\sigma}_{\theta/2}\bm{\nabla}$ (2)
and the non-Hermitian bottom layer Hamiltonian is
$h_{b}=-i\hbar v_{F}\bm{\sigma}_{-\theta/2}\bm{\nabla}+i\lambda_{V}\sigma_{z}.$ (3)
Here, $i\lambda_{V}\sigma_{z}$ is a balanced gain and loss imposing on the bottom layer, which induces non-Hermiticity, and $\bm{\sigma}_{\theta/2}$ corresponds to the rotated Pauli matrices $e^{-i\theta\sigma_{z}/4}(\pm\sigma_{x},\sigma_{y})e^{i\theta\sigma_{z}/4}$, with $\pm$ for $K$ and $K^{\prime}$ valley, respectively. $\textbf{T}(\textbf{r})=\sum^{3}_{j=1}T_{j}e^{-ik_{\theta}\textbf{q}_{j}\cdot\textbf{r}}$ is the sublattice-dependent moiré potential that couples the two layers. The interlayer coupling is in the form
$T_{j+1}=w_{0}\sigma_{0}+w_{1}(\sigma_{x}\cos j\phi+\sigma_{y}\sin j\phi)$ (4)
with $\phi=2\pi/3$ and $w_{0}(w_{1})$ being the interlayer coupling strength in the AA (AB) region. The unit vectors are $\textbf{q}_{1}=(0,-1)$ and $\textbf{q}_{2,3}=(\pm\sqrt{3}/2,1/2)$. $k_{\theta}=2k_{D}\sin(\theta/2)$ is the moiré modulation vector and $k_{D}=4\pi/(3a_{0})$ is the magnitude of the Dirac wave vector, where $a_{0}$ is the lattice constant of monolayer graphene. The hopping strengths $w_{0}$ and $w_{1}$ encode the inter-layer couplings in the AA and AB regions respectively. Correspondingly, dimensionless parameters can be defined as $\alpha_{i}=\frac{w_{i}}{\hbar v_{F}k_{\theta}}$ for $i=0,1$.
@@
If there is no AA stacking ($w_{0}=0$), the Hamiltonian preserves chiral symmetry $\mathcal{C}\mathcal{H}(\textbf{r})\mathcal{C}^{\dagger}=-\mathcal{H}(\textbf{r})$ under the chiral symmetry operator $\mathcal{C}=\sigma_{z}$ in the absence of non-Hermiticity [50, 51, 49]. In the presence of the balanced gain and loss $i\sigma_{z}$ term, the chiral symmetry is broken but the non-Hermitian chiral symmetry is preserved, which is crucial for the presence of non-Hermitian topology.
In the following numerical calculations, we take $w_{1}$ = 110 meV, $\hbar v_{F}/a_{0}=2.413eV$, and $w_{0}=0$ in the chiral limit. In the absence of non-Hermiticity, the first magic angle is given by $\alpha_{1H}\approx 0.593$, which corresponds to $\theta\approx 1.05$.
III Exceptional rings
Figure 2: (Color online) In the presence of $\lambda_{V}$, exceptional rings appear. (a) The real part and (b) the imaginary part of the lowest-energy moiré band spectrum. The exceptional rings are marked by red line segments. The twist angle is taken as $\alpha_{1}=0.2$ and the non-Hermitian strength is $\lambda_{V}=0.07$ eV.
@@ III Exceptional rings
Let us investigate the band structure for a twist angle away from the Hermitian magic angle. In the absence of non-Hermiticity, this is not a magic angle with a flat band. If the interlayer coupling effect is absent ($w_{0}=w_{1}=0$), the dispersion in the mBZ is equivalent to the folded dispersion in the Brillouin zone of graphene. Since the physics in the $K$ valley and in the $K^{\prime}$ valley are related, we only discuss the $K$ valley in the following. The Dirac point at the $K$ corner in the top layer is mapped to $K_{M}$ point in the mBZ, while that at the $K$ corner of the bottom layer is mapped to $K^{\prime}_{M}$ point in the mBZ.
@@
In the presence of the non-Hermitian perturbation $\lambda_{V}$, the energy dispersion becomes complex. In Fig. 2(a), we can see that the non-Hermitian perturbation deforms the Dirac cones at $K^{\prime}_{M}$ corners. More specifically, the degeneracy point at the $K^{\prime}_{M}$ points in the real part of the spectrum morphs into a ring of degeneracy points. This is a stark difference from single-layer graphene with on-site gain and loss. The latter has deformed Dirac cones in both $K$ and $K^{\prime}$ corners in the Brillouin zone. This difference is due to the fact that the Dirac cones at the $K_{M}$ and $K_{M}^{\prime}$ points in non-Hermitian TBG are folded from different graphene layers. The non-Hermitian $\lambda_{V}$ can only deform the Dirac cone on the bottom layer.
@@
In particular, the degeneracy of the two lowest-energy states leads to the appearance of singularity points in momentum space. At the same time, a ring of degeneracy points appears at the same momentum position in the imaginary part of the spectrum, as shown in Fig. 2(b). We found that not only do the eigenvalues coalesce but also their corresponding eigenvectors coalesce at this point. This confirms that these singularity points are exceptional points in momentum space. Thus, the Dirac cone is turned into a ring of exceptional points.
Note that the spawning of exceptional ring out of a Dirac cone in non-Hermitian TBG is similar to that in other two-dimensional non-Hermitian semimetals protected by non-Hermitian chiral symmetry [52, 53, 54, 55]. It is worthwhile to investigate the symmetry in non-Hermitian TBG.
IV Tripod model and non-Hermitian chiral symmetry